• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Bosted

    Log entry time 08:06:53 on November 6,2002

    Entry number 89727

    keyword=Shift Summary

    Ready to start Q3-g50 kinamtics, but there are problems at injector at beginning
    of shift, so we wait for beam.

    At 01:30, MCC called to say problem is fixed, and they will start to tune for
    100 muA beam. Started optics target run at 1:37.

    2:00 Finished optics run. Switched to 15 cm LD2. This took 15 minutes (10
    minutes for MCC to mask the beam). Started production running.

    During physics replay of first production run (3024), noticed timing cuts were
    off. Changed them to 170 to 180 nsec (like Q3D-g40) to match data, then
    found 7 good coincidences for 1M triggers, which seems about right.

    At 4:40 MCC called to say they will take a q.e. measurement for 5 minutes

    At 5:30 had minor coda problem (no ROCS responding), so run 3027 was ended
    early. Restarting a new run fixed the problem

    Beam was good until 7:40, when injector problems developed, and MCC took
    the beam for tuning